OpenStudy (anonymous):

(-1/4)(x-8)+(1/2)(x+2)=x+9 please show me how to solve this

OpenStudy (anonymous):

distribute first to get (-1/4)x+2 + (1/2)x+1 = x +9

OpenStudy (anonymous):

next, collect like terms: (-1/4)x+(1/2)x + 2 +1 = x + 9 (1/4)x+3 = x +9

OpenStudy (anonymous):

bring x's to one side and numbers to the other x-(1/4)x = 3-9 (3/4)x = -6

OpenStudy (anonymous):

multiply by the reciprocal of (3/4) on both sides to get x = -6(4/3) = -24/3 = -8

OpenStudy (anonymous):

how did you get neg 8 from 6(4/3) to me it would be -7 1/3 I see how you mulitiplied -6 by 4 but y four and not 3 3/3 is the whole number not 4 im confused

OpenStudy (anonymous):

-6 times 4/3 is the same as -6/1 times 4/3. Multiply the numerators and denominators separately in fractions to get (-6)(4)/(1)(3) = -24/3 = -8

OpenStudy (anonymous):

Ok i get it thank you

OpenStudy (anonymous):

No problem